Roybus, Inc., a manufacturer of flash memory, just reported that its main production facility in Taiwan was destroyed in a fire. While the plant was fully insured, the loss of production will decrease Roybus's free cash flow by $183 million at the end of this year and by $63 million at the end of next year.

a. If Roybus has 37 million shares outstanding and a weighted average cost of capital of 13.5%, what change in Roybus's stock price would you expect upon this announcement? (Assume that the value of Roybus' debt is not affected by the event.)

b. Would you expect to be able to sell Roybus's stock on hearing this announcement and make a profit? Explain.

Having trouble working the problem I know its going to look something like this: 183/(1+0.135)-63(1+0.135)^2 For question A, with a negative number? please provide steps.
Answer #1

a.

change in value = -FCF1/(1+WACC)-FCF2/(1+WACC)^2

=-183/(1+0.135)-63/(1+0.135)^2

=-210.1379

price drop = change in value/shares outstanding=(-210.1379)/37=-5.68

b.

No, markets are extremely efficient, and unless you are a full time trader it will be difficult to react quickly and capitalize on the news